A container 80cm/60cm/70cm needs to be filled up completely with water. Tube A supplies 24l of water per minute and Tube B supplies 18l of water per minute. Mr Tan fills up the container by taking turns to use both Tubes for 17minutes. How long does he use each of the Tubes?
Found 3 solutions by mananth, ikleyn, n2:
Answer by mananth(16949)

Let tube A be open for x minutes
Tube B for y minutes
x+y=17........................(1)

The capacity of tank is 336 liters (70x60x80)/1000
Pipe B will fill the tank in 336/24=14 minutes
Pipe A will fill the tank in 336/18=18.66 minutes
18.66x+14y=336...............................(2)

x= Time open for A
y= Time open for B

1 x + 1 y = 17 .............1
Tank filling by both tubes
19 x + 24 y = 336 .............2
Eliminate y
multiply (1)by -24
Multiply (2) by 1
-24 x -24 y = -408
19 x + 24 y = 336
Add the two equations
-5 x = -72
/ -5
x = 13
plug value of x in (1)
1 x + 1 y = 17
13 + y = 17
y = 17 -13.48
y = 4
y = 4
x= 13 Time open for A
y= 4 Time open for B
m.ananth@hotmail.ca

Answer by ikleyn(53996)

.
A container 80cm/60cm/70cm needs to be filled up completely with water.
Tube A supplies 24l of water per minute and Tube B supplies 18l of water per minute.
Mr Tan fills up the container by taking turns to use both Tubes for 17minutes.
How long does he use each of the Tubes?
~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution in the post by @mananth is  FATALLY  INCORRECT.

        It is incorrect,  because his setup equations are irrelevant to the problem.

        I came to bring a correct solution.


The volume of the tank is 80*60*70 = 336000 cm^3,  or 336 liters.

Let x be the time for tube A working, in minutes;

then (17-x) is the time for the tube B working.


Write an equation for the total volume

    24*x + 18*(17-x) = 336.


Simplify and find x

    24x + 306 - 18x = 336

    24x - 18x = 336 - 306

        6x    =    30

         x    =    30/6 = 5.


So, tube A worked 5 minutes;  tube B worked 17-5 = 12 minutes.    <<<---===  ANSWER


CHECK.  The total volume is  24*5 + 18*12 = 336 liters,  which is precisely correct.

Solved correctly.

This situation is  100%  typical.

@mananth uses a computer code,  which generates solutions for his problems.
If a problem is one step from a standard,  the code produces gibberish,
but @mananth's reaction is zero - he  NEVER  reads what his code does produce:
he submits it to the forum without reading it,  without checking and without looking at it.

It is why every his post should be checked by others and it is why I check as many of them as I can.



Answer by n2(95)

.
A container 80cm/60cm/70cm needs to be filled up completely with water.
Tube A supplies 24l of water per minute and Tube B supplies 18l of water per minute.
Mr Tan fills up the container by taking turns to use both Tubes for 17minutes.
How long does he use each of the Tubes?
~~~~~~~~~~~~~~~~~~~~~~~~~~~


The volume of the tank is 80*60*70 = 336000 cm^3,  or 336 liters.

Let x be the time for tube A working, in minutes;

then (17-x) is the time for the tube B working.


Write an equation for the total volume

    24*x + 18*(17-x) = 336.


Simplify and find x

    24x + 306 - 18x = 336

    24x - 18x = 336 - 306

        6x    =    30

         x    =    30/6 = 5.


So, tube A worked 5 minutes;  tube B worked 17-5 = 12 minutes.    <<<---===  ANSWER


CHECK.  The total volume is  24*5 + 18*12 = 336 liters,  which is precisely correct.

Solved.